OpenStudy (anonymous):

help meh! Assume that the value of an automobile is inversely proportional to its age. If it is worth $3,000 when it is two years old, what is its value when it is six years old?

OpenStudy (anonymous):

inverse variation?

OpenStudy (diyadiya):

\[Value=k \frac{1}{age}\]

OpenStudy (anonymous):

as age increases the value decreases, it is inversely proportional. Older -> less valuable (<3000)

OpenStudy (diyadiya):

\[3000=k*\frac{1}{2}\] Now find k and find the value when age is 6

OpenStudy (anonymous):

like 500?

OpenStudy (diyadiya):

What did you get k as?

OpenStudy (diyadiya):

\[3000=k*\frac{1}{2} \ (or) \ 3000= \frac{k}{2}\] Now calculate k

OpenStudy (anonymous):

am I supposed to put the 6 over it?

OpenStudy (diyadiya):

No now dont use 6 Just calculate k first

OpenStudy (anonymous):

how ?!

OpenStudy (diyadiya):

\[3000=k*\frac{1}{2} \ (or) \ 3000= \frac{k}{2} \] Did you understand this?

OpenStudy (anonymous):

oh wait... 6,000...I'm stupid

OpenStudy (diyadiya):

Yup Right :) Now you found k Lets Go back to the equation \[Value=k \frac{1}{age}\]

OpenStudy (diyadiya):

Substitute k=6000 and age=6 and find the value :)

OpenStudy (anonymous):

1,000?!

OpenStudy (diyadiya):

Yup :)

OpenStudy (anonymous):

thank you so much, diya! Sorry I'm a retard...

OpenStudy (diyadiya):

Lol No problem :) i'm not a nerd either :P